# Consensus Planning with Primal, Dual, and Proximal Agents

### v2

@@ 4.1 Overview of Results
We prove in this section different forms of convergence of Algorithm 1 (3ACP). We saw in Section 3.3 that when the agents are exclusively primal, dual, or proximal, the algorithm reduces to Bregman ADMM, dual ascent, and regular ADMM, respectively. Individually, each of these algorithms is known to converge sublinearly when the functions are merely assumed to be generally convex, and strongly convex in the case of dual ascent. Under stricter conditions, namely strong-convexity and Lipschitz continuous gradients, they reach linear convergence rates (see [DY16, HL17] for ADMM and [WB14, LLF22] for Bregman ADMM).
@@ 4.2 Preliminaries and Assumptions
Let $\mathcal{M}=\mathcal{P}\cup\mathcal{D}\cup\mathcal{X}$ be a set of primal ($\mathcal{P}$), dual ($\mathcal{D}$), and proximal agents ($\mathcal{X}$), with objective functions $g_{i},\leavevmode\nobreak\ i\in\mathcal{M}$. For all $i\in\mathcal{M}$, the functions $g_{i}$ are convex, and in particular $\mu_{i}$-strongly convex for $i\in\mathcal{D}$, and with $\beta_{i}$-Lipschitz continuous gradients for $i\in\mathcal{P}$.
These assumptions are the weakest we can require, and the ones we will use to prove the $O\left(\frac{1}{k}\right)$ convergence of the algorithm. The strong convexity of the dual agents is necessary for the convergence of dual ascent, which is a special case of 3ACP, while the Lipschitz-continuity of the primal agents is necessary in order to bound them above by a quadratic function.
@@
$\mathcal{L}(\mathbf{x}^{*},z^{*},\bm{\lambda})\leq\mathcal{L}(\mathbf{x}^{*},z^{*},\bm{\lambda}^{*})\leq\mathcal{L}(\mathbf{x},z,\bm{\lambda}^{*}),\qquad\forall\leavevmode\nobreak\ \mathbf{x},z,\bm{\lambda}.$ (10)
@@
To prove linear convergence results, as well as to tighten the convergence bounds even in the sublinear case, we will assume that all agents have strongly convex objective functions with Lipschitz continuous gradients.
Assumption 4.4.
Let $\mathcal{M}$ be a set of primal, dual, and proximal agents, with objective functions $g_{i},\leavevmode\nobreak\ i\in\mathcal{M}$. For all $i\in\mathcal{M}$, the functions $g_{i}$ are $\mu_{i}$-strongly convex with $\beta_{i}$-Lipschitz continuous gradients.
Remark 4.5.
  • When $\mathcal{D}=\emptyset$, $r^{k}$ can also be expressed as:
  • When $\mathcal{M}=\mathcal{D}$, we have:
    $V^{k}$ $=\sum_{i\in\mathcal{D}}D_{\psi_{i}}(\lambda_{i}^{k},\lambda_{i}^{*}),$
    $r^{k}$ $=\sum_{i\in\mathcal{D}}D_{\psi_{i}}(\lambda_{i}^{k},\lambda_{i}^{k-1}),$
@@ 4.3 Plain Convergence
Consider the consensus problem (1) and suppose Assumptions  4.1 and 4.2 hold. Then, the iterates generated by Algorithm 1 satisfy the following inequality for all $k\geq 0$:
@@
$\sum_{i\in\mathcal{M}}\frac{\mu_{i}}{2}\lVert x_{i}^{k+1}-x_{i}^{*}\rVert^{2}\leq V^{k}-V^{k+1}-r^{k+1},\leavevmode\nobreak\ \text{and }\sum_{i\in\mathcal{D}}\frac{\mu_{i}}{2L_{i}^{2}}\lVert\lambda_{i}^{k}-\lambda_{i}^{*}\rVert^{2}\leq V^{k}-V^{k+1}-r^{k+1}.$ (12)
@@
Consider the consensus problem (1) and suppose Assumptions  4.1, 4.2 , and 4.3 hold. Then, for the iterates generated by Algorithm 1 we have:
@@
$z^{k+1}-x_{i}^{k+1}\to 0,\leavevmode\nobreak\ \forall i\in\mathcal{M}.$
@@
$\lambda_{i}^{k}$ $\to\lambda_{i}^{*},\leavevmode\nobreak\ \forall i\in\mathcal{D},$
$x_{i}^{k}$ $\to x_{i}^{*},\leavevmode\nobreak\ \forall i\in\mathcal{M},$
@@ 4.5 Linear Convergence
We prove in this section the two-step linear convergence of the 3ACP Algorithm under the stronger assumptions that all agents’ functions be strongly convex with Lipschitz continuous gradients (Assumption 4.4) rather than the milder assumptions that only the dual agents be strongly convex, and only the primal agents have Lipschitz continuous gradients. We first define a few terms to simplify the notation, letting $\mathcal{S}\subset\mathcal{M}$ be any subset of agents:
@@
Note that if all the agents are of the same type, we recover known results about the (one-step) linear convergence of the algorithm, corresponding to the linear convergence of ADMM in the case of all proximal agents (see [LLF22, Thm 3.4]), of Bregman ADMM in the case of all primal agents (see [LLF22, Thm 3.8]), and of dual ascent in the case of all dual agents. These results are summarized below.
@@ 5.1 Acceleration
The algorithms considered so far are essentially first-order methods, only making use of gradient information in some form or another. Such algorithms can often be accelerated through schemes that use previous iterates in the updates. One important such acceleration scheme is Nesterov’s accelerated gradient descent [Nes83], which has been successfully ported to many first order algorithms, including ADMM. This is especially interesting in our context since the primal and dual agents are handled as approximations of proximal agents, and are thus likely to lose some performance with respect to the latter.
Given that ADMM works both on the primal (individual plans and consensus) and dual (prices) variables, acceleration techniques may target either of those sets of variables, or both. For example, [GOSB14] presents an accelerated ADMM where the second set of primal variable and the prices are accelerated, while [KCSB15] only modifies the dual variable but requires an additional update of the primal variables. [OHTG13] considers accelerating the linearized ADMM algorithm used for the primal agents and is thus particularly relevant to us. Additionally, accelerated methods are known to not be monotone in the objective value, and they usually display rippling effects. Adaptive restart strategies to alleviate these issue were proposed in [OC15], and they can be adapted to the ADMM acceleration schemes, as was the case in [GOSB14].
One difficulty is combining these different schemes. We can readily apply the accelerated scheme of [GOSB14] when the agents are dual and/or proximal, while we can also directly apply the algorithm of [OHTG13] when the agents are all primal, and some results resulting from those implementations are presented in Section 6. Combining these in order to allow for the acceleration of the algorithm for any combination of agents will be an interesting research topic.
@@ 5.2 Tighter Quadratic Bounds for Primal Agents
Our approach to primal agents involved the use of a quadratic approximation of the objective functions $g_{i}$ of the form $x\mapsto g_{i}(x_{i}^{k})+\nabla g_{i}(x_{i}^{k})^{T}(x-x_{i}^{k})+\frac{L_{i}}{2}\lVert x\rVert^{2}$. This nonetheless restricts the type of linearization to spherical quadratic functions. We can tighten the bounds by considering not just spherical quadratic functions but more general quadratic functions so long as they dominate $g_{i}$. Letting $H_{i}$ be a symmetric positive definite matrix such that $H_{i}\succeq\nabla^{2}g_{i}(x),\leavevmode\nobreak\ \forall x$, we may then use the following function $\phi_{i}$ in the definition of the Bregman divergence $D_{\phi_{i}}$: $\phi_{i}(x)=\frac{1}{2}x^{T}H_{i}x-g_{i}(x)=\frac{1}{2}\lVert x\rVert_{H_{i}}^{2}-g_{i}(x)$, instead of $\frac{L_{i}}{2}x^{T}x-g_{i}(x)$.
@@ 5.3 Second Order Information
Related to Section 5.2 above, we might actually have access to second order information. This is especially useful for the primal and dual agents, since in the former case we could make use of a (close to) second-order linearization, while in the latter case, the price update could result from a Newton (or quasi-Newton) update as opposed to a simple first-order dual ascent.
For example, for primal agents, we could consider Bregman divergence $D_{\phi_{i}}$ letting $\phi_{i}(x)=\frac{1}{2}x^{T}(\nabla^{2}g_{i}(x)+(\beta_{i}-\|\nabla^{2}g_{i}(x)\|)I)x-g_{i}(x)$ or $\phi_{i}(x)=\frac{1}{2}x^{T}(\nabla^{2}g_{i}(x)+(\beta_{i}-\mu_{i})I)x-g_{i}(x)$.
For dual agents, we could consider a price update of the form $\lambda_{i}^{k+1}=\lambda_{i}^{k}+\rho_{i}\nabla^{2}g_{i}(x_{i}^{k+1})(z^{k+1}-x_{i}^{k+1})$, or alternatively $\lambda_{i}^{k+1}=\lambda_{i}^{k}+\rho_{i}(\nabla^{2}g_{i}(x_{i}^{k+1})+\epsilon I)(z^{k+1}-x_{i}^{k+1})$, for some $\epsilon>0$, where $\rho_{i}$ and $\epsilon$ are appropriately chosen to guarantee convergence. In that case, however, the consensus update would need to be modified to:
@@ 5.4 Choice of Hyperparameters
In the absence of knowledge of $\mu_{i}$ and/or $\beta_{i}$, we can conceive of evaluating them “on the fly” based on accumulated information, or alternatively, dynamically adjusting them, as suggested by [HYW00]. In particular, the self-adaptive parameter tuning can be applied at an agent level by considering their respective primal and and dual residuals.
@@
5.5 Regularized Dual Updates
One scheme we considered, but discarded for now as it didn’t seem to have much effect on numerical examples, was to update the dual agents through a two step procedure, as follows:
$\tilde{x}_{i}^{k+1}$ $=\arg\min_{x}g_{i}(x)-{\lambda_{i}^{k}}^{T}x,$
$x_{i}^{k+1}$ $=\frac{L_{i}\tilde{x}_{i}^{k+1}+\rho_{i}z^{k}}{L_{i}+\rho_{i}},$
for some $L_{i}>0$, although we do not require that $L_{i}\geq\beta_{i}$ unlike in the primal case. $x_{i}^{k+1}$ is thus written as a weighted average of $\tilde{x}_{i}^{k+1}$ (which would have been the regular update in dual ascent) and $z_{k}$. The plan is thus regularized towards the consensus. This nonetheless put restrictions on the admissible values of $(\rho_{i},L_{i})$. This procedure can be motivated through several observations:
  • The optimality condition for the problem yielding $\tilde{x}_{i}^{k+1}$ is $\nabla g_{i}(\tilde{x}_{i}^{k+1})=\lambda_{i}^{k}$. As a result, $x_{i}^{k+1}$ can be expressed as:
    $x_{i}^{k+1}$ $=\arg\min_{x}g_{i}(\tilde{x}_{i}^{k+1})+\nabla g_{i}(\tilde{x}_{i}^{k+1})^{T}x+\frac{L_{i}}{2}\lVert x-\tilde{x}_{i}^{k+1}\rVert^{2}-{\lambda_{i}^{k}}^{T}x+\frac{\rho_{i}}{2}\lVert z^{k}-x\rVert^{2}.$
    This is an update of the same form as the linearized ADMM update (3) for the primal agent, where the linearization is here performed at $\tilde{x}_{i}^{k+1}$, but because we do not require that $L_{i}\geq\beta_{i}$, the linearization does not necessarily yield an upper bound of $g_{i}$.
  • Suppose that all the agents are dual agents and we use $\rho$ as a learning rate for the price update and a common value $L=L_{i},\forall i$. Performing dual ascent using this two step approach yields the following steps:
    $\tilde{x}_{i}^{k+1}$ $=\arg\min_{x_{i}}g(x_{i})-{\lambda_{i}^{k}}^{T}x_{i},$
    $x_{i}^{k+1}$ $=\frac{L\tilde{x}_{i}^{k+1}+\rho_{i}z^{k}}{L+\rho_{i}},$
    $z^{k+1}$ $=\frac{1}{|\mathcal{D}|}\sum_{i}x_{i}^{k+1}=\frac{L\tilde{z}^{k+1}+\rho_{i}z^{k}}{L+\rho},$
    $\lambda_{i}^{k+1}$ $=\lambda_{i}^{k}+\rho_{i}(z^{k+1}-x_{i}^{k+1})=\lambda_{i}^{k}+\frac{\rho_{i}L}{\rho_{i}+L}\left(\tilde{z}^{k+1}-\tilde{x}_{i}^{k+1}\right).$
    We observe that this would result in updates that are the ones of a dual ascent procedure with rate $\alpha=\frac{\rho L}{\rho+L}$, which also imposes the condition $\rho(L-\mu)\leq\mu L$ since the dual ascent update requires that $\alpha\leq\mu$.
  • Suppose the function $g_{i}$ is quadratic of the form $g_{i}(x)=\frac{1}{2}Lx^{T}x+b^{T}x$. Solving the problem as a proximal agent problem yields $x_{i}^{k+1}=\frac{\lambda_{i}^{k}-b+\rho z^{k}}{L+\rho}$. On the other hand, using the above two step procedure leads to $\tilde{x}_{i}^{k+1}=\frac{\lambda_{i}^{k}-b}{L}$, and then $x_{i}^{k+1}=\frac{L\tilde{x}_{i}^{k+1}+\rho z^{k}}{L+\rho}=\frac{\lambda_{i}^{k}-b+\rho z^{k}}{L+\rho}$, which is the same update as the one for the proximal agent. As a result, in the case of a spherical quadratic function, any dual agent can be turned into a proximal agent through this regularization. Solving the dual agent problem and then regularizing is equivalent to directly solving the proximal agent problem.
@@ 6.1 Agents
    This would also be the exact same form of the updates for dual agents, were we to use the regularized dual update suggested in Section 5.5.
@@ 8 Acknowledgments
This paper is part of a stream of work to develop a market-based approach to coordination, the Consensus Planning Protocol (CPP) [DPJvR22, vR24]. We would especially like to thank Garrett van Ryzin for his support, as well as Manan Chopra, Shekhar Jain, Mitchell Joblin, and Allam Reddy for their feedback.
@@ Appendix A Proofs
$\nabla g_{i}(x_{i}^{k+1})$ $=\lambda_{i}^{k}+\rho_{i}(z^{k}-x_{i}^{k+1})-\left(\nabla\phi_{i}(x_{i}^{k+1})-\nabla\phi_{i}(x_{i}^{k})\right),\leavevmode\nobreak\ i\in\mathcal{P}$ (13)
$\nabla g_{i}(x_{i}^{k+1})$ $=\lambda_{i}^{k}+\rho_{i}(z^{k}-x_{i}^{k+1}),\leavevmode\nobreak\ i\in\mathcal{X},$ (14)
$\nabla g_{i}(x_{i}^{k+1})$ $=\lambda_{i}^{k},\leavevmode\nobreak\ i\in\mathcal{D}.$ (15)
@@
00 0 $\leq\mathcal{L}(\mathbf{x}^{k+1},z^{k+1},\bm{\lambda}^{*})-\mathcal{L}(\mathbf{x}^{*},z^{*},\bm{\lambda}^{*})$
@@
00 0 $\leq\mathcal{L}(\mathbf{x}^{k+1},z^{k+1},\bm{\lambda}^{*})-\mathcal{L}(\mathbf{x}^{*},z^{*},\bm{\lambda}^{*})$
@@
00 0 $\leq\mathcal{L}(\mathbf{x}^{k+1},z^{k+1},\bm{\lambda}^{*})-\mathcal{L}(\mathbf{x}^{*},z^{*},\bm{\lambda}^{*})$
@@
00 0 $\leq\mathcal{L}(\mathbf{x}^{k+1},z^{k+1},\bm{\lambda}^{*})-\mathcal{L}(\mathbf{x}^{*},z^{*},\bm{\lambda}^{*})\leq V^{k}-V^{k+1}-r^{k+1},$
@@
    $\leavevmode\nobreak\ \qquad+\sum_{i\in\mathcal{P}}\left(\nabla\phi_{i}(x_{i}^{k+1})-\nabla\phi_{i}(x_{i}^{k})\right)^{T}(x_{i}^{k+1}-x_{i}^{*})$
@@
The application of gradient Lipschitz continuity to Equation (11) would have added a term $-\frac{1}{2L_{i}}\lVert\nabla g_{i}(x_{i}^{k+1})-\nabla g_{i}(x_{i}^{*})\rVert^{2}$ to the right of Equation (16) in the proof of Proposition 4.6, and yielded:
@@
$\tilde{\rho}_{i}$ =$=\begin{cases}\rho_{i}&i\in\mathcal{P}\cup\mathcal{X},\\ 0&i\in\mathcal{D}\end{cases},$ $\phi_{i}(x)$ $=0,\leavevmode\nobreak\ i\in\mathcal{D}\cup\mathcal{X}.$
@@
  • Recall that since $-g_{i}^{*}$ has $\frac{1}{\mu_{i}}$-Lipschitz continuous gradients, we have according to Lemma B.3 that $D_{-g_{i}^{*}}(\lambda_{i}^{k+1},\lambda_{i}^{*})\leq\frac{1}{2\mu_{i}}\lVert\lambda_{i}^{k+1}-\lambda_{i}^{*}\rVert^{2}$, and so, using (24) once again:

### v3

@@ 4.1 Overview of Results
We prove in this section different forms of convergence of Algorithm 1 (3ACP). We saw in Section 3.3 that when the agents are exclusively primal, dual, or proximal, the algorithm reduces to Bregman ADMM, dual ascent, and regular ADMM, respectively. Individually, each of these algorithms is known to converge sublinearly when the functions are merely assumed to be generally convex, and strongly convex in the case of dual ascent. Under stricter conditions, namely strong-convexity and Lipschitz continuous gradients, they reach linear convergence rates (see deng2016global; hong2017linear for ADMM and wang2014bregman; lin2022alternating for Bregman ADMM).
@@ 4.2 Preliminaries and Assumptions
Let $\mathcal{M}=\mathcal{P}\cup\mathcal{D}\cup\mathcal{X}$ be a set of primal ($\mathcal{P}$), dual ($\mathcal{D}$), and proximal agents ($\mathcal{X}$), with objective functions $g_{i},\penalty 10000\ i\in\mathcal{M}$. For all $i\in\mathcal{M}$, the functions $g_{i}$ are convex, and in particular $\mu_{i}$-strongly convex for $i\in\mathcal{D}$, and with $\beta_{i}$-Lipschitz continuous gradients for $i\in\mathcal{P}$.
These assumptions are to be expected: the strong convexity of the dual agents is necessary for the convergence of dual ascent, which is a special case of 3ACP, while the Lipschitz-continuity of the primal agents is necessary in order to bound them above by a quadratic function.
@@
$\mathcal{L}(\mathbf{x}^{*},z^{*},\bm{\lambda})\leq\mathcal{L}(\mathbf{x}^{*},z^{*},\bm{\lambda}^{*})\leq\mathcal{L}(\mathbf{x},z,\bm{\lambda}^{*}),\qquad\forall\penalty 10000\ \mathbf{x},z,\bm{\lambda}.$ (10)
@@ 4.3 Plain Convergence
Consider the consensus problem (1) and suppose Assumptions 4.1 and 4.2 hold. Then, the iterates generated by Algorithm 1 satisfy the following inequality for all $k\geq 0$:
@@
$\sum_{i\in\mathcal{M}}\frac{\mu_{i}}{2}\lVert x_{i}^{k+1}-x_{i}^{*}\rVert^{2}\leq V^{k}-V^{k+1}-r^{k+1},\penalty 10000\ \text{and }\sum_{i\in\mathcal{D}}\frac{\mu_{i}}{2L_{i}^{2}}\lVert\lambda_{i}^{k}-\lambda_{i}^{*}\rVert^{2}\leq V^{k}-V^{k+1}-r^{k+1}.$ (12)
@@
Consider the consensus problem (1) and suppose Assumptions 4.1, 4.2 , and 4.3 hold. Then, for the iterates generated by Algorithm 1 we have:
@@
$z^{k+1}-x_{i}^{k+1}\to 0,\penalty 10000\ \forall i\in\mathcal{M}.$
@@
$\lambda_{i}^{k}$ $\to\lambda_{i}^{*},\penalty 10000\ \forall i\in\mathcal{D},$
$x_{i}^{k}$ $\to x_{i}^{*},\penalty 10000\ \forall i\in\mathcal{M},$
@@ 4.5 Linear Convergence
We prove in this section the two-step linear convergence of the 3ACP Algorithm. To do so, as well as to tighten the convergence bounds even in the sublinear case, we will assume that all agents have strongly convex objective functions with Lipschitz continuous gradients, rather than the milder assumptions that only the dual agents be strongly convex, and only the primal agents have Lipschitz continuous gradients.
Assumption 4.9.
Let $\mathcal{M}$ be a set of primal, dual, and proximal agents, with objective functions $g_{i},\penalty 10000\ i\in\mathcal{M}$. For all $i\in\mathcal{M}$, the functions $g_{i}$ are $\mu_{i}$-strongly convex with $\beta_{i}$-Lipschitz continuous gradients.
Remark 4.10.
  • When $\mathcal{D}=\emptyset$, $r^{k}$ can also be expressed as:
  • When $\mathcal{M}=\mathcal{D}$, we have:
    $V^{k}$ $=\sum_{i\in\mathcal{D}}D_{\psi_{i}}(\lambda_{i}^{k},\lambda_{i}^{*}),$
    $r^{k}$ $=\sum_{i\in\mathcal{D}}D_{\psi_{i}}(\lambda_{i}^{k},\lambda_{i}^{k-1}),$
We then first define a few terms to simplify the notation, letting $\mathcal{S}\subset\mathcal{M}$ be any subset of agents:
@@
We are only able to achieve a “two-step” linear convergence here (a linear relation between values two steps apart, $V^{k}$ and $V^{k+2}$), as opposed to the more common one-step linear convergence (relating consecutive values $V^{k}$ and $V^{k+1}$), which is the one found in the special cases where all the agents are of the same type. This can be understood by the slight shift in discretization implied by plain gradient schemes compared to proximal algorithms. Methods such as gradient descent can be interpreted as forward Euler methods, while proximal algorithms have backward Euler method interpretation (parikh2014proximal). To reconcile these two forms of approximation, we need to consider the progress after two steps of the algorithm.
Note that if all the agents are of the same type, we recover known results about the (one-step) linear convergence of the algorithm, corresponding to the linear convergence of ADMM in the case of all proximal agents (see (lin2022alternating, Thm 3.4)), of Bregman ADMM in the case of all primal agents (see (lin2022alternating, Thm 3.8)), and of dual ascent in the case of all dual agents. These results are summarized below.
@@ 5.1 Acceleration
The algorithms considered so far are essentially first-order methods, only making use of gradient information. Such algorithms can often be accelerated through schemes that use previous iterates in to modify the direction of the step update, and use momentum. This is different from the ergodic convergence analysis of Section 4.4, where we considered the convergence of the running average of the iterates, but where that did not modify the step updates. One important such acceleration scheme is Nesterov’s accelerated gradient descent (nesterov1983method), which has been successfully ported to many first order algorithms, including ADMM. This is especially interesting in our context since the primal and dual agents are handled as approximations of proximal agents, and are thus likely to lose some performance with respect to the latter.
Given that ADMM works both on the primal (individual plans and consensus) and dual (prices) variables, acceleration techniques may target either of those sets of variables, or both. For example, goldstein2014fast presents an accelerated ADMM where the second set of primal variable and the prices are accelerated, while kadkhodaie2015accelerated only modifies the dual variable but requires an additional update of the primal variables. ouyang2013stochastic considers accelerating the linearized ADMM algorithm used for the primal agents and is thus particularly relevant to us. Additionally, accelerated methods are known to not be monotone in the objective value, and they usually display rippling effects. Adaptive restart strategies to alleviate these issue were proposed in o2015adaptive, and they can be adapted to the ADMM acceleration schemes, as was the case in goldstein2014fast.
One difficulty is combining these different schemes. We can readily apply the accelerated scheme of goldstein2014fast when the agents are dual and/or proximal, while we can also directly apply the algorithm of ouyang2013stochastic when the agents are all primal, and some results resulting from those implementations are presented in Section 6. Combining these in order to allow for the acceleration of the algorithm for any combination of agents will be an interesting research topic.
@@ 5.2 Tighter Quadratic Bounds for Primal Agents
Our approach to primal agents involved the use of a quadratic approximation of the objective functions $g_{i}$ of the form $x\mapsto g_{i}(x_{i}^{k})+\nabla g_{i}(x_{i}^{k})^{T}(x-x_{i}^{k})+\frac{L_{i}}{2}\lVert x\rVert^{2}$. This nonetheless restricts the type of linearization to spherical quadratic functions. We can tighten the bounds by considering not just spherical quadratic functions but more general quadratic functions so long as they dominate $g_{i}$. Letting $H_{i}$ be a symmetric positive definite matrix such that $H_{i}\succeq\nabla^{2}g_{i}(x),\penalty 10000\ \forall x$, we may then use the following function $\phi_{i}$ in the definition of the Bregman divergence $D_{\phi_{i}}$: $\phi_{i}(x)=\frac{1}{2}x^{T}H_{i}x-g_{i}(x)=\frac{1}{2}\lVert x\rVert_{H_{i}}^{2}-g_{i}(x)$, instead of $\frac{L_{i}}{2}x^{T}x-g_{i}(x)$.
@@ 5.3 Second Order Information
Related to Section 5.2 above, we might actually have access to second order information. This is especially useful for the primal and dual agents, since in the former case we could make use of a (close to) second-order approximation, while in the latter case, the price update could result from a Newton (or quasi-Newton) update as opposed to a simple first-order dual ascent. The use of second-order information in decentralized or federated learning has been considered in works such as zhang2015disco or wang2018giant for example.
For example, for primal agents, we could consider a variable Bregman divergence $D_{\phi_{i}^{k}}$ of the form $\phi_{i}^{k}(x)=\frac{1}{2a_{i}}x^{T}(\nabla^{2}g_{i}(x_{i}^{k})+\epsilon I)x-g_{i}(x)$, where $\varepsilon>0$ is used to ensure strong convexity. If we assume that the functions $g_{i}$ are strongly convex, this would yield updates similar to the Newton-type updates of the FedHybrid algorithm (niu2023fedhybrid), reinforcing the view of FedHybrid as a particular case of Bregman ADMM.
For dual agents, we could consider a price update of the form $\lambda_{i}^{k+1}=\lambda_{i}^{k}+\rho_{i}\nabla^{2}g_{i}(x_{i}^{k+1})(z^{k+1}-x_{i}^{k+1})$, or alternatively $\lambda_{i}^{k+1}=\lambda_{i}^{k}+\rho_{i}(\nabla^{2}g_{i}(x_{i}^{k+1})+\epsilon I)(z^{k+1}-x_{i}^{k+1})$, for some $\epsilon>0$, where $\rho_{i}$ and $\epsilon$ are appropriately chosen to guarantee convergence, in particular so that $\rho_{i}(\nabla^{2}g_{i}(x_{i}^{k+1})+\epsilon I)||<\mu_{i}$. In that case, however, the consensus update would need to be modified to:
@@ 5.4 Choice of Hyperparameters
In the absence of knowledge of $\mu_{i}$ and/or $\beta_{i}$, we can conceive of evaluating them “on the fly” based on accumulated information, or alternatively, dynamically adjusting them, as suggested by he2000alternating. In particular, the self-adaptive parameter tuning can be applied at an agent level by considering their respective primal and and dual residuals.
@@ 8 Acknowledgments
This paper is part of a stream of work to develop a market-based approach to coordination, the Consensus Planning Protocol (CPP) (cpp; cpptalk). We would especially like to thank Garrett van Ryzin for his support, as well as Manan Chopra, Shekhar Jain, Mitchell Joblin, and Allam Reddy for their feedback.
@@ Appendix A Proofs
$\nabla g_{i}(x_{i}^{k+1})$ $=\lambda_{i}^{k}+\rho_{i}(z^{k}-x_{i}^{k+1})-\left(\nabla\phi_{i}(x_{i}^{k+1})-\nabla\phi_{i}(x_{i}^{k})\right),\penalty 10000\ i\in\mathcal{P}$ (13)
$\nabla g_{i}(x_{i}^{k+1})$ $=\lambda_{i}^{k}+\rho_{i}(z^{k}-x_{i}^{k+1}),\penalty 10000\ i\in\mathcal{X},$ (14)
$\nabla g_{i}(x_{i}^{k+1})$ $=\lambda_{i}^{k},\penalty 10000\ i\in\mathcal{D}.$ (15)
@@
$0$ $\leq\mathcal{L}(\mathbf{x}^{k+1},z^{k+1},\bm{\lambda}^{*})-\mathcal{L}(\mathbf{x}^{*},z^{*},\bm{\lambda}^{*})$
@@
$0$ $\leq\mathcal{L}(\mathbf{x}^{k+1},z^{k+1},\bm{\lambda}^{*})-\mathcal{L}(\mathbf{x}^{*},z^{*},\bm{\lambda}^{*})$
@@
$0$ $\leq\mathcal{L}(\mathbf{x}^{k+1},z^{k+1},\bm{\lambda}^{*})-\mathcal{L}(\mathbf{x}^{*},z^{*},\bm{\lambda}^{*})$
@@
$0$ $\leq\mathcal{L}(\mathbf{x}^{k+1},z^{k+1},\bm{\lambda}^{*})-\mathcal{L}(\mathbf{x}^{*},z^{*},\bm{\lambda}^{*})\leq V^{k}-V^{k+1}-r^{k+1},$
@@
    $\penalty 10000\ \qquad+\sum_{i\in\mathcal{P}}\left(\nabla\phi_{i}(x_{i}^{k+1})-\nabla\phi_{i}(x_{i}^{k})\right)^{T}(x_{i}^{k+1}-x_{i}^{*})$
@@
The application of gradient Lipschitz continuity to Equation (11) would have added a term $\frac{1}{2L_{i}}\lVert\nabla g_{i}(x_{i}^{k+1})-\nabla g_{i}(x_{i}^{*})\rVert^{2}$ to the right of Equation (16) in the proof of Proposition 4.4, and yielded:
@@
$\tilde{\rho}_{i}$ $=\begin{cases}\rho_{i}&i\in\mathcal{P}\cup\mathcal{X},\\ 0&i\in\mathcal{D}\end{cases},$ $\phi_{i}(x)$ $=0,\penalty 10000\ i\in\mathcal{D}\cup\mathcal{X}.$
@@
  • Recall that since $g_{i}^{*}$ has $\frac{1}{\mu_{i}}$-Lipschitz continuous gradients, we have according to Lemma B.3 that $D_{-g_{i}^{*}}(\lambda_{i}^{k+1},\lambda_{i}^{*})\leq\frac{1}{2\mu_{i}}\lVert\lambda_{i}^{k+1}-\lambda_{i}^{*}\rVert^{2}$, and so, using (24) once again: